# A Survey on Vision-Language-Action Models for Embodied AI

### v2

@@ III-B8 LLM-based Control Policies
Figure 5: Different approaches to connect LLM to multi-modal modules in high-level task planner.
III-C High-level Task Planner
A high-level task planner aims to divide a complex task into a series of executable sub-tasks. This approach is often referred to as task decomposition, and is also related to the classical Task and Motion Planning (TAMP) setting. Many high-level task planners are constructed on top of LLMs. While integrating multi-modal modules into LLMs in an end-to-end fashion (§III-C1) is a common approach, training with multi-modal data can be expensive. Consequently, some task planners explore to utilize language (§III-C3) or code (§III-C4) as the medium for exchanging multi-modal information instead, as they can be natively processed by LLMs.
III-C1 End-to-end
While control policies are effective in understanding and executing simple language instructions, they tend to struggle in long-horizon tasks that involve multiple sub-tasks. Large language models have been identified as powerful tools to interpret such complex long-horizon tasks. Consequently, several approaches are dedicated to integrating LLMs as high-level task planners. Their goal is to decompose long-horizon tasks into simpler sub-tasks, enabling low-level control policies to execute them sequentially, contributing to the development of a hierarchical robotic system. Table IV outlines important details about these high-level task planners.
@@ III-C1 End-to-end
Table IV: High-level task planners. VL: vision-language fusion. Sim/Real: simulated/real-world. Mani/Navi: manipulation/navigation.
Model V-Encoder Lang Model VL Control Policy Train Robotic Tasks
SayCan [183] CNN PaLM [185]
FLAN [186]
Concat BC-Z [130]
MT-Opt [12]
Yes Real-Navi&Mani (Everyday Robots): kitchen
LID [184] CNN GPT-2 [187] Concat - Yes Sim: VirtualHome
[…]
Codex [190]
- - No Sim: VirtualHome
(SL)3 [191] ResNet18 [29] T5-small [10] Xattn HLSM [192] Yes Sim-Navi: ALFRED
EmbodiedGPT [101] EVA ViT-G/14 [193] LLaMA-7B [150] Embodied-former MLP Yes Sim-Mani: Franka Kitchen, Meta-World
PaLM-E [181] ViT [30],
OSRT [194]
PaLM [185] Concat Interactive Language policy [117], RT-1 [3] Yes Sim: TAMP; Sim&Real-Mani: Lang.-Table; Real: SayCan setup
LEO [195] ConvNeXt [196], PointNet++ [197] Vicuna-7B [125] Concat MLP Yes Sim-Mani: CLIPort task subset; Sim-Navi: Habitat-web
3D-LLM [198] CLIP, 3D-CLR [199], ConceptFusion [200] Flamingo [11], BLIP-2 [16] Concat DD-PPO [201] Yes Sim-Navi: Habitat
StructFormer [202], StructDiffusion [203] PCT [204] - Concat [205], RRT-connect [206] Yes Sim-Mani: PyBullet; Real-Mani (Franka): rearrange
ShapeLLM [207] ReCon++ LLaMA Concat - Yes 3D MM-Vet Benchmark
Inner Monologue [208] MDETR [156], CLIP InstructGPT [209]
PaLM [185]
Language CLIPort, MDETR, SayCan policies No Sim&Real-Mani: Tabletop Rearrangement; Real-Navi&Mani: Kitchen Mobile Mani.
LLM-Planner [210] HLSM object detector [192] GPT-3 [189] Language HLSM [192] Yes Sim-Navi: ALFRED
SMs [211] CLIP, ViLD [155] GPT-3 [189], RoBERTa [212] Language & Code CLIPort [13] No Sim-Mani: pick, place
ProgPrompt [213] ViLD [155] GPT-3 [189] Code API [214, 215] No Sim: VirtualHome
ChatGPT for Robotics [216] YOLOv8 [217] ChatGPT [1] Code API No Real-Navi: drone flight; Sim-Navi: AirSim, Habitat
CaP [218] ViLD [155], MDETR [156] GPT-3 [189]
Codex [190]
Code API No Sim-Mani & Sim-Navi: pick, place, etc.
DEPS [219] CLIP ChatGPT [220] Code MC-Controller, Steve-1 No Sim: Minecraft
ConceptGraphs [221] SAM [31], CLIP, LLaVA [45] GPT-4 Code (JSON) API No Sim: AI2-THOR; Real (Spot Arm): pick, place
[…]
Semi-Supervised Skill Learning with Latent Language ((SL)3) [191] is a learning algorithm that alternates between three steps: segmentation, labeling, and parameter update. In the segmentation step, high-level subtasks are aligned with low-level actions, subtask descriptions are then inferred in the labeling step, and finally, the network parameters are updated. This approach enables a hierarchical policy to discover reusable skills with sparse natural language annotations.
PaLM-E [181] integrates ViT [30] and PaLM [185] to create a large embodied multi-modal language model capable of performing high-level embodied reasoning tasks. Based on perceived images and high-level language instructions, PaLM-E generates a text plan that serves as instructions for low-level robotic policies. In a mobile manipulation environment, they map the generated plan to executable low-level instructions with SayCan [183]. As the low-level policy executes actions, PaLM-E can also replan based on changes in the environment. With PaLM as its backbone, PaLM-E can handle both normal visual question answering (VQA) tasks, along with the additional embodied VQA tasks.
III-C2 3D Vision
3D information is vital in embodied tasks, and thus it is also explored for task planners. Because the majority of current VLMs deal with images as visual inputs, it requires architectural change to VLMs to incorporate 3D visual inputs and therefore they are usually end-to-end models.
LEO [195] identifies the conventional use of image inputs to be a limiting factor for multi-modal generalist agents in interacting with the 3D world. This novel approach involves training an LLM-based architecture using a new dataset with two stages. The first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning. LEO demonstrates proficiency not only in 3D captioning and question-answering tasks but also in embodied tasks, including embodied reasoning, embodied navigation, and robotic manipulation.
StructFormer [202] addresses the task of arranging objects into complex structures. Previous methods rely on pairwise semantic relations and sequential manipulation, while StructFormer explicitly reasons about the relationships among all objects and predicts their target positions. StructDiffusion [203] extends this by incorporating a diffusion-based generation approach.
Table V: Recent datasets for robot learning. Following the definition in RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects. $\ast$ Dataset collected in simulators rather than in the real world. Some datasets are continuously updated; this table is based on the information from their original papers. This table is partially based on [224, 149].
@@
MIME [225] Human Teleop RGB, D Demo Baxter 12 20 1 - 8.3K
RoboTurk [226] $\ast$ Human Teleop RGB - Sawyer 2 - 1 - 2.1K
@@
BC-Z [130] Human Teleop RGB Lang, Demo Everyday 3 100 1 - 25.9K
RT-1 [3] Human Teleop RGB Lang Everyday 12 700+ 2 16 130K
MOO [129] Human Teleop RGB Multi-modal Everyday 5 - - 106 59.1K
@@
RH20T [224] Human Teleop RGB, D Lang (4 robots) 42 147 7 - 110K+
DROID [149] Human Teleop RGB, D Lang Franka Panda 86 - 564 - 76K
@@
III-C3 Language-based
Inner Monologue [208] sits between high-level command and low-level policy to enable closed-loop control planning. It employs LLM to generate language instructions for low-level control policies and dynamically updates these instructions based on feedback received from the control policies. The feedback encompasses various sources: success feedback, object and scene feedback, and human feedback. As the feedback is communicated to the LLM in textual format, no additional training is required for the LLM. A similar approach is also proposed in ReAct [104], which supports more flexible reasoning traces, allowing for more diverse types of reasoning for various tasks.
III-C4 Code-based
ProgPrompt [213] introduces a novel task-planning approach by prompting LLMs with program-like specifications detailing available actions and objects. This enables LLMs to generate high-level plans for household tasks in a few-shot manner. Environmental feedback can be incorporated through assertions within the program. This prompting scheme leverages the world knowledge and programming skills of LLMs.
ChatGPT for Robotics [216] take advantage of the programming ability of ChatGPT to facilitate “user on the loop” control, a departure from the conventional “engineer in the loop” methodology. The procedure includes several steps: firstly, a list of APIs is defined, such as an object-detection API, a grasp API, a move API; secondly, a prompt is then constructed for ChatGPT, specifying the environment, API functionality, task goal, etc.; thirdly, iteratively prompting ChatGPT to write code with the defined APIs that can execute the task, provided the access to simulation and user feedback for evaluating the code quality and safety; finally, executing the ChatGPT generated code. In this procedure, ChatGPT serves as a high-level task planner, similar to PaLM-E [181], and actions are generated through function calls to corresponding low-level APIs.
Code as policies (CaP) [218] also leverages the code-writing capability of LLMs. It employs GPT-3 [189] or Codex [190] to generate policy code, which, in turn, invokes perception modules and control APIs. CaP exhibits proficiency on spatial-geometric reasoning, generalization to new instructions, and parameterization for low-level control primitives. By leveraging the multimodal capabilities of GPT-4V, COME-robot [230] eliminates the need for perception APIs in CaP. This also opens up possibilities for open-ended reasoning and adaptive planning within a closed-loop framework, enabling capabilities such as failure recovery and free-form instruction following.
@@ III-C4 Code-based
Pros and Cons. While end-to-end task planners like SayCan [183] share a similar architecture to low-level control policies and can be optimized for specific tasks, their training costs can be prohibitive due to the large model size of the combined LLM and visual transformer. Language-based task planners offer the advantage of seamless integration with existing language-conditioned control policies. However, they often require finetuning or alignment methods to map the generated plan to executable language instructions for low-level control policies. On the other hand, code-based task planners leverage the programming capability of LLMs to connect perception and action modules. This approach does not require additional training, but its performance may be constrained by the capabilities of existing models.
IV Datasets and Benchmarks
Table VI: Simulators for benchmarking VLA models. D: Depth. Seg: segmentation. A: Audio. N: Normals. Force: simulated contact force between the end-effector and the item. PD: Pre-defined. Vers: versions. This table is partially based on [231, 232].
@@ IV Datasets and Benchmarks
UI Engine Task Observation Action Agent Description Related
Gibson [233] 572/- - - Pybullet Navi RGB, D, N, Seg - - Navi only -
iGibson [234, 235, 232] 15/108 152/5 Mouse, VR Pybullet Navi, Mani RGB, D, Seg, N, Flow, LiDAR Force Husky, TurtleBot v2, LoCoBot, etc. VR, Continuous Extended States. Vers: iGibson 0.5, 1.0, 2.0 Benchmarks: BEHAVIOR-100 [231], BEHAVIOR-1K [236]
SAPIEN [237] - 2346/- Code PhysX Navi, Mani RGB, D, Seg Force Franka Articulation, Ray Tracing VoxPoser
AI2-THOR [238] -/120 118/118 Mouse Unity Navi, Mani RGB, D, Seg, A Force, PD ManipulaTHOR, LoCoBot, etc. Object States, Task Planning. Vers: [239, 240] Benchmarks: ALFRED, RoPOR [241]
VirtualHome [242] 7/- -/509 Lang Unity Navi, Mani RGB, D, Seg, F Force, PD Human Object States, Task Planning LID, Translated〈LM〉, ProgPrompt
TDW [243] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, Seg, A Force Fetch, Sawyer, Baxter Audio, Fluids -
RLBench [119] 1/- 28/28 Code Bullet Mani RGB, D, Seg Force Franka Tiered Task Difficulty Hiveformer, PerAct
@@
Habitat [247, 248] Matterport + Gibson Mouse Bullet Navi RGB, D, Seg, A Force Fetch, Franka, AlienGO Fast, Navi only. Vers: Rearrangement [249], Habitat 2.0 VC-1, PACT; OVMM [250]
ALFRED [251] -/120 84/84 - Unity Navi, Mani RGB, D, Seg PD Human Diverse long-horizon tasks (SL)3, LLM-Planner
@@
The collection of real-world robot data presents significant challenges due to various factors. Firstly, the collection process is hindered by the substantial costs associated with procuring robot equipment, setting up environments, and dedicating extensive human resources. Secondly, the gathering of expert demonstration data requires considerable time investments. Thirdly, the diverse types and configurations of robots introduce inconsistency in sensory data, control modes, gripper types, etc. Lastly, achieving precision in capturing object 6D poses and exactly replicating or resetting setups remains elusive. As a result of these difficulties, public real-world robot datasets are relatively scarce. Furthermore, evaluating a robot system’s performance in real-world conditions introduces another layer of complexity, as it is challenging to precisely reproduce a setup and often requires human supervision. We summarize the robot datasets from recent VLAs in Table V.
@@
Consequently, many researchers resort to simulated environments as a means to mitigate these obstacles and accelerate the data collection process. We compare the most relevant simulators in Table VI. Nevertheless, this strategy presents its own challenges, chief among them being the sim-to-real gap. This discrepancy arises when models trained on simulated data exhibit poor performance in real-world deployment. The causes of this gap are multifaceted, encompassing differences in rendering quality, inaccuracies in physics simulations, and domain shifts characterized by unrealistic object properties and robot motion planners. For instance, simulating non-rigid objects like liquids or towels presents significant difficulties. Moreover, incorporating new objects into simulators requires considerable effort, often involving techniques such as 3D scanning and mesh editing. Despite these hurdles, simulated environments provide automated evaluation metrics that aid researchers in consistently evaluating robotic models. Many benchmarks are based on simulators because only simulated environments can reproduce the experimental setup exactly, while real-world evaluation varies between different models, making it infeasible to compare them reliably. The HomeRobot OVMM Benchmark [250] proposes a sim-to-real benchmark, yet its consistency remains to be observed.
@@
Several approaches advocate for automated data collection. RoboGen [254] employs a generative simulation paradigm that proposes interesting skills, simulates corresponding environments, and selects optimal learning approaches to train policies for acquiring those skills. AutoRT [255] functions as a robot orchestrator driven by LLMs, generating tasks, filtering them by affordance, and utilizing either autonomous policies or human teleoperators to collect and evaluate data. DIAL [256] focuses on augmenting language instructions in existing datasets using VLMs. RoboPoint [124] generates scenes procedurally with randomized 3D layouts, objects, and camera viewpoints.
@@
An alternative strategy to address the data scarcity issue in real-world settings is to leverage human data. Human behavior offers plentiful guidance for robot policies due to its dexterity and diversity. However, this strategy also comes with inherent drawbacks. Capturing and transferring human hand/body motions to robot embodiments is inherently difficult. Moreover, the inconsistency in human data poses a hurdle, as some data may be egocentric while others are captured from third-person perspectives. Additionally, filtering human data to extract useful information can be labor-intensive. These obstacles underscore the complexities involved in incorporating human data into robot learning processes. Due to limited space, we refer interested readers to a comprehensive comparison of existing datasets in [257].

### v4

@@ IV-A3 End-to-end Task Planners with 3D Vision
Figure 6: Different approaches to connect LLM to multi-modal modules in modular task planners.
IV-B Modular Task Planners
Finetuning end-to-end models on embodied data can be expensive, and there are approaches that adopt a modular design by assembling off-the-shelf LLMs and VLMs into task planners.
IV-B1 Language-based Task Planners
Language-based approaches use natural language descriptions as the medium for exchanging multimodal information, as illustrated in Figure 6a.
Inner Monologue [9] sits between high-level command and low-level policy to enable closed-loop control planning. It employs LLM to generate language instructions for low-level control policies and dynamically updates these instructions based on feedback received from the control policies. The feedback encompasses various sources: success feedback, object and scene feedback, and human feedback. As the feedback is communicated to the LLM in textual format, no additional training is required for the LLM. A similar approach is used in ReAct [92].
@@ IV-B1 Language-based Task Planners
Table V: Recent datasets for robot learning. Following the definition in RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects. $\ast$ Dataset collected in simulators rather than in the real world. Some datasets are continuously updated; this table is based on the information from their original papers. Table adapted from [180, 146].
@@
MIME [181] Human Teleoperation RGB, D Demo Baxter 12 20 1 - 8.3K
RoboTurk [182] $\ast$ Human Teleoperation RGB - Sawyer 2 - 1 - 2.1K
@@
BC-Z [113] Human Teleoperation RGB Lang, Demo Everyday 3 100 1 - 25.9K
RT-1-Kitchen [114] Human Teleoperation RGB Lang Everyday 12 700+ 2 16 130K
MOO [111] Human Teleoperation RGB Multi-modal Everyday 5 - - 106 59.1K
@@
RH20T [180] Human Teleoperation RGB, D Lang (4 robots) 42 147 7 - 110K+
DROID [146] Human Teleoperation RGB, D Lang Franka Panda 86 - 564 - 76K
@@
IV-B2 Code-based Task Planners
Code-based task planners leverage the coding ability of LLMs to generate task plans in the form of a program. Object detectors, VLMs, and control policies can be invoked via APIs, as shown in Figure 6b.
ProgPrompt [171] introduces a novel task-planning approach by prompting LLMs with program-like specifications detailing available actions and objects. This enables LLMs to generate high-level plans for household tasks in a few-shot manner. Environmental feedback can be incorporated through assertions within the program.
ChatGPT for Robotics [172] take advantage of the programming ability of ChatGPT to facilitate “user on the loop” control, a departure from the conventional “engineer in the loop” methodology. The procedure includes several steps: firstly, a list of APIs is defined, such as an object-detection API, a grasp API, a move API; secondly, a prompt is then constructed for ChatGPT, specifying the environment, API functionality, task goal, etc.; thirdly, iteratively prompting ChatGPT to write code with the defined APIs that can execute the task, provided the access to simulation and user feedback for evaluating the code quality and safety; finally, executing the ChatGPT generated code. In this procedure, ChatGPT serves as a high-level task planner and actions are generated through function calls to corresponding low-level APIs.
Code as policies (CaP) [174] also leverages the code-writing capability of LLMs. It employs GPT-3 or Codex to generate policy code, which, in turn, invokes perception modules and control APIs. CaP exhibits proficiency on spatial-geometric reasoning, generalization to new instructions, and parameterization for low-level control primitives. By leveraging the multimodal capabilities of GPT-4V, COME-robot [187] eliminates the need for perception APIs in CaP. This also opens up possibilities for open-ended reasoning and adaptive planning within a closed-loop framework, enabling capabilities such as failure recovery and free-form instruction following.
@@ IV-B2 Code-based Task Planners
Pros and Cons. Monolithic task planners that utilize grounded task planning focus on generating executable plans. End-to-end models like PaLM-E share an architecture similar to most LVLAs and can be finetuned on specialized embodied data to achieve better performance. However, the training costs of such large models can be a significant concern. In contrast, modular task planners are more readily deployable because they leverage off-the-shelf LLMs and VLMs. Language-based task planners offer the advantage of seamless integration of LLMs and VLMs, as they are designed to operate in the natural language space. However, they often require extra steps to align the generated task plans with language instructions that are admissible to low-level control policies. On the other hand, while code-based task planners may require manually wrapping VLMs and control policies in APIs and preparing clear documentation in advance, they enable code debugging and provide greater controllability. Nevertheless, their performance can be constrained by the programming capabilities of existing models.
V Datasets and Benchmarks
Table VI: Simulators and simulated benchmarks. Control: Continuous control tasks. D, S, A, N: Depth, segmentation, audio, normals. Force: simulated contact force between end-effector and item. PD: Pre-defined. Table adapted from [188, 189].
@@ V Datasets and Benchmarks
UI Physics Engine Task Observation Action Agent Description Related
Gibson [190] 572/- - - Pybullet Navi RGB, D, N, S - - Navi only -
iGibson [191, 192, 189] 15/108 152/5 Mouse, VR Pybullet Navi, Mani RGB, D, S, N, Flow, LiDAR Force TurtleBot v2, LoCoBot, etc. VR, Continuous Extended States. Versions: iGibson 0.5, 1.0, 2.0 Benchmarks: BEHAVIOR-100 [188], BEHAVIOR-1K [193]
SAPIEN [194] - 2346/- Code PhysX Navi, Mani RGB, D, S Force Franka Articulation, Ray Tracing VoxPoser. Benchmarks: SIMPLER [195]
AI2-THOR [196] -/120 118/118 Mouse Unity Navi, Mani RGB, D, S, A Force, PD ManipulaTHOR, LoCoBot, etc. Object States, Task Planning. Versions: [197, 198] Benchmarks: ALFRED, RoPOR [199]
VirtualHome [200] 7/- -/509 Lang Unity Navi, Mani RGB, D, S Force, PD Human Object States, Task Planning LID, Translated〈LM〉, ProgPrompt
TDW [201] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, S, A Force Fetch, Sawyer, Baxter Audio, Fluids -
RLBench [102] 1/- 28/28 Code Bullet Mani RGB, D, S Force Franka Tiered Task Difficulty Hiveformer, PerAct
@@
Habitat [205, 206] Matterport + Gibson Mouse Bullet Navi RGB, D, S, A Force Fetch, Franka, AlienGO Fast, Navi only. Versions: Rearrangement [207], Habitat 2.0 VC-1, PACT; OVMM [208]
ALFRED [209] -/120 84/84 - Unity Navi, Mani RGB, D, S PD Human Diverse long-horizon tasks (SL)3, LLM-Planner
@@
Genesis [212] (Rigid, deformable, liquid, etc.) Code (Propri-etary) Navi, Mani RGB, D, S, N Force Franka, Unitree, etc. High-speed comprehensive physics simulation -
V-A Real-world Robot Datasets & Benchmarks
Embodied AI faces significant data scarcity issues because real-world robot data is not as readily available as language data. Collecting real-world robot datasets poses multiple challenges. Firstly, it is impeded by the cost and time required to procure robotic equipment, set up environments, and gather expert data through dedicated policies or human teleoperation. Secondly, the diverse types and configurations of robots introduce inconsistencies in sensory data, control modes, gripper types, etc. Lastly, accurately capturing object 6D poses and reproducing setups remains elusive. We summarize recent robot datasets in Table V. In addition, real-world benchmarks are further complicated by the need for human evaluation.
V-B Simulators, Simulated Robot Datasets & Benchmarks
@@
Many researchers resort to simulated environments to circumvent real-world obstacles and scale the data collection process. We compare simulators and simulated benchmarks in Table VI. Nevertheless, this strategy presents its own challenges, chief among which is the sim-to-real gap. This discrepancy arises when models trained on simulated data exhibit poor performance during real-world deployment. The causes of this gap are multifaceted, encompassing unrealistic rendering quality, inaccuracies in physics simulations, and domain shifts characterized by object properties and robot motion planners. For instance, simulating non-rigid objects such as deformable objects or liquids presents significant difficulties. Moreover, importing new objects into simulators requires considerable effort, often involving techniques such as 3D scanning and mesh editing. Despite these hurdles, simulated environments provide automated evaluation metrics that aid researchers in consistently evaluating robotic models. Many benchmarks are based on simulators because they can precisely reproduce the experimental setup and yield fair comparisons of different models. Another technique, known as real-to-sim, can improve simulation fidelity, recreate failure cases, or facilitate digital twins.
V-C Automated Dataset Collection
@@
Several approaches advocate for automated dataset collection. RoboGen [213] employs a generative simulation paradigm that proposes interesting skills, simulates corresponding environments, and selects optimal learning approaches to train policies for acquiring those skills. AutoRT [214] functions as a robot orchestrator driven by LLMs, generating tasks, filtering them by affordance, and utilizing either autonomous policies or human teleoperators to collect and evaluate data. DIAL [215] focuses on augmenting language instructions in existing datasets using VLMs. RoboPoint [107] generates scenes procedurally with randomized 3D layouts, objects, and camera viewpoints.
V-D Human Datasets
@@
An alternative strategy to address data scarcity in real-world settings is to leverage human data. Human behavior offers plentiful guidance for robot policies due to its dexterity and diversity. However, this strategy also comes with inherent drawbacks. Capturing and transferring human hand/body motions to robot embodiments is inherently difficult. Moreover, the inconsistency in human data poses a hurdle, as some data may be egocentric while others are captured from third-person perspectives. Additionally, filtering human data to extract useful information can be labor-intensive. These obstacles underscore the complexities involved in incorporating human data into robot learning processes. UMI [216] proposes a method to mitigate these issues using hand-held grippers. For a more comprehensive comparison of human datasets, we refer interested readers to [217].
V-E Task Planning Benchmarks
EgoPlan-Bench [218] focuses on benchmarking real-world task planning with human annotations. PlanBench [219, 220] comprehensively assesses various aspects of task planning ability, such as cost optimality, plan verification, and replanning. LoTa-Bench [221] directly evaluates task planning by executing the generated plans in simulators and calculating success rates. Embodied Agent Interface (EAI) [222] argues that this approach fails to pinpoint issues in LLMs. By formalizing the input-output of LLM-based modules for decision-making tasks, EAI enables more fine-grained metrics beyond success rates.
V-F Embodied Question Answering Benchmarks
Embodied question answering (EQA) benchmarks, as summarized in Table VII, do not directly evaluate robotic tasks like manipulation and navigation, but they are aimed at other relevant abilities for embodied AI, such as spatial reasoning, physics understanding, as well as world knowledge. EQA is akin to previous visual question answering benchmarks, but differs in that the agent can actively explore the environment before providing an answer. EmbodiedQA [223] and IQUAD [224] were among the first works to introduce this type of benchmarks. MT-EQA [225] focuses on complex questions involving multiple targets. MP3D-EQA [226] converts previous RGB inputs to point clouds, testing 3D perception capabilities.
Active exploration requires access to a simulator, limiting the types of data that can be used, such as real-world videos. EgoVQA [227] shifts the focus of VQA to egocentric videos. EgoTaskQA [228] emphasizes spatial, temporal, and causal relationship reasoning. EQA-MX [229] investigates multimodal expressions (MX), including regular verbal utterances and nonverbal gestures like eye gaze and pointing. OpenEQA [230] evaluates seven main categories, including functional reasoning and world knowledge.